OpenStudy (anonymous):

If f(x) = x +2, all over 2, then what is f^–1(x)

OpenStudy (anonymous):

Assuming this an inverse function, the first step you would take is switching x and ys, then solving for y again.

OpenStudy (mathteacher1729):

First write \[y = \dfrac{x+2}{2}\] Now swap the x and y. \[x = \dfrac{y+2}{2}\] Now solve for y

OpenStudy (mathteacher1729):

The reason we "swap the x and y" is because the inverse of a function is a reflection about the line y = x. ( so wherever we see y, we switch it with x and vice versa.)

OpenStudy (anonymous):

so you get 2y = x?

OpenStudy (anonymous):

2y-2 = x*